The number of distinct solutions of sin 5x * cos 3x = sin 9x * cos 7x in [0,π/2] is

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

2 sin 5x cos 3x = 2 sin 9x cos 7x
Sin 8x + sin 2x = sin 16x + sin 2x

Sin 8x = 2 sin 8x cos 8x

Cos 8x = 1/2

8x = pi/3
x = pi/24

Regards